OpenStudy (anonymous):

2^a=5^b=20^c.Express c in term of a and b.

OpenStudy (anonymous):

Let 2^a=5^b=20^c=m nw 2=m^1/a 5=m^1/b 20=m^1/c we know that 2.2.5=20 therefore we can write m^1/a X m^1/a X m^1/b=m^1/c =>1/a+1/a+1/b=1/c =>2/a+1/b=1/c

OpenStudy (anonymous):

Thank for reply and answer.